# Segment Any Change

### v1

@@ 5 Conclusion
Appendix A Pseudocode of Bitemporal Latent Matching
Appendix B Implementation Details
B.1 Baselines for zero-shot change proposal and detection
SAM forward: For object proposal generation, we adopt a point per side of 64, an NMS threshold of 0.7, a predicted IoU threshold of 0.5, and a stability score threshold is 0.8 for LEVIR-CD, S2Looking, SECOND, and 0.95 for xView2. To obtain the image embedding with a constant norm (), we demodulate the output of the image encoder of SAM with the affine transformation () of the last layer normalization, i.e., , where is the single input image.
@@ B.1 Baselines for zero-shot change proposal and detection
AnyChange: For a fair comparison and automatic benchmark evaluation, we use fully automatic mode for AnyChange in Table 1. The change angle threshold (155) is obtained by OTSU and a linear search on the small validation set sampled from the SECOND training set. This threshold is directly used for the other three datasets without any adjustment.
@@
B.2 Pseudo-label Training

### v4

@@ 5 Conclusion
Acknowledgements
This work was supported in part by ARO (W911NF-21-1-0125), ONR (N00014-23-1-2159), the CZ Biohub, and the National Natural Science Foundation of China under Grant No. 42325105.
Appendix A Appendix / supplemental material
Appendix B Pseudocode of Bitemporal Latent Matching
Appendix C Implementation Details
C.1 Baselines for zero-shot change proposal and detection
SAM forward: For object proposal generation, we adopt a point per side of 64, an NMS threshold of 0.7, a predicted IoU threshold of 0.5, and a stability score threshold is 0.8 for LEVIR-CD, S2Looking, SECOND, and 0.95 for xView2. To obtain the image embedding with a constant norm (), we demodulate the output of the image encoder of SAM with the affine transformation () of the last layer normalization, i.e., , where is the single input image. The forward computation was conducted on 8 NVIDIA A4000 GPUs.
@@ C.1 Baselines for zero-shot change proposal and detection
AnyChange: For a fair comparison and automatic benchmark evaluation, we use fully automatic mode for AnyChange in Table 1. The change angle threshold (155∘) is obtained by OTSU and a linear search on the small validation set sampled from the SECOND training set. This threshold is directly used for the other three datasets without any adjustment.
@@
C.2 Pseudo-label Training
@@
Appendix D Limitations
Zero-shot change detection is an open problem in remote sensing and computer vision communities. There is little reference to point out any potential effective roadmap before our work. Our work is the first to define this problem suitably and provide a simple yet effective model and evaluation protocol. The problem formulation and evaluation protocol themselves bring some potential limitations (i.e., scenario coverage, the robustness to objects of different geometries) since there is little mature infrastructure, e.g., a concept-complete class-agnostic change detection dataset.
Appendix E Broader Impacts
The proposed method can detect class-agnostic changes in Earth’s surface, however, its actual effect is impacted by SAM’s latent space. The model may produce some impossible changes due to SAM’s biases. These issues warrant further research and consideration when building upon this work for real-world applications.
Appendix F More visualization
Nature Image Domain. Our method can be also used in the natural image domain to detect more general object changes, as shown in Fig. 5.
More demonstration of point query mechanism. We provide an additional example to supplement Fig. 4, which includes unchanged and changed buildings simultaneously. It is more clear to show that AnyChange can more accurately detect building changes with the help of the point query.
Effectiveness of AnyChange in detecting tiny object changes. Fig. 7 demonstrates a case of tiny/minor changes, such as small vehicle changes. We observe that directly applying AnyChange to the original image overlooks these subtle changes (see the first row). After we bilinearly upsampled the red box region by 2 and then applied AnyChange to it, we find some tiny/minor changes could be detected (see the second row). This observation shows that our method has the ability on tiny object change detection, although it is not yet optimal. Future work could use our approach as a strong baseline to further improve the detection of subtle changes.